Does principle of superposition also hold for non-linear PDEs? If not, then give an example of non-linear PDE where principle of superposition does not hold.
1
Expert's answer
2020-09-06T18:41:34-0400

Unlike linear PDEs, with non-linear PDEs the principle of superposition (which states that linear combination of solutions of equation is again the solution of this equation) does not hold in general case. Example:

"u_{x} +u^2 u_y=0" (1)

One solutions of this PDE is "u_1 = \frac{-1+\sqrt{1+4xy}}{2x}" . However, the function "u =cu_1" does not solve the same PDE unless "c = 0, \pm1".